OpenStudy (anonymous):

Solve :5n+1 = 3n-5 ____ ____ 8 4

OpenStudy (ranga):

Just multiply both sides by 8, bring all n terms to one side and constants to the other side and solve for n.

OpenStudy (anonymous):

\[\frac{5n+1}{8}=\frac{3n-5}{4}\]Multiply both sides by 8.\[\frac{8}{8}(5n+1)=\frac{8}{4}(3n-5)\]Evaluate the fractions.\[1(5n+1)=2(3n-5)\]Distribute.\[5n+1=6n-10\]Subtract 6n from both sides.\[-n+1=-10\]Subtract 1 from both sides.\[-n=-11\]Multiply both sides by -1.\[n=11\]